# Lotus at SemEval-2025 Task 11: RoBERTa with Llama-3 Generated Explanations for Multi-Label Emotion Classification

### v2

@@ 3.3 Challenges and Solutions
Emotion Text + Explanation (RoBERTa) Text Only (RoBERTa) Text Only (Llama-3)
Precision Recall F1 Precision Recall F1 Precision Recall F1
Anger 0.6695 0.6304 0.6479 0.6892 0.5116 0.5871 0.7337 0.4193 0.5336
Fear 0.7983 0.8739 0.8343 0.8009 0.8200 0.8149 0.7658 0.8387 0.8006
Joy 0.7957 0.7291 0.7581 0.7587 0.6925 0.7232 0.7971 0.6567 0.7201
Sadness 0.6831 0.8127 0.7423 0.7636 0.6935 0.7268 0.6624 0.7662 0.7105
Surprise 0.7625 0.6702 0.7132 0.7248 0.6877 0.7039 0.6091 0.6008 0.6049
@@ 4 Experimental Setup
We evaluated our multi-label emotion detection approach, using RoBERTa fine-tuned with explanatory content generated by Llama-3, on the BRIGHTER dataset.
Text preprocessing and tokenization were performed with the RobertaTokenizer from Hugging Face. In the first phase, Llama-3 generated explanations, which were concatenated with the original text. In the second phase, both the original text and the explanation were tokenized together to help the model learn the emotional context.
For fine-tuning Llama-3, we used 4-bit quantization and LoRA, with a batch size of 2, gradient accumulation steps of 4, and a learning rate of $1\times 10^{-4}$ for 30 training steps. These explanations were used in the second phase for emotion classification. RoBERTa was fine-tuned in the second phase with binary emotion labels, using a batch size of 8, a learning rate of $5\times 10^{-5}$, and 3 epochs. Model performance was evaluated using precision, recall, and F1-scores, including both Macro and Micro F1-scores to assess multi-label classification.
@@ 5 Results
In this section, we present the performance of our system, Lotus, on the competition task. Using the Text + Explanation (RoBERTa) method, Lotus scored 0.7319, outperforming the SemEval Baseline (0.7083) but falling short of the top score of 0.823. Ranked 44th, it performed competitively, though there is room for improvement to reach the highest positions.
@@ 5.1 Performance Comparison for Individual Emotions
In Table 2, the performance of Lotus was compared across individual emotions using three methods: Text + Explanation (RoBERTa), Text Only (RoBERTa), and Text Only (Llama-3), focusing on precision, recall, and F1-score.
Metric Text + Explanation (RoBERTa) Text Only(RoBERTa) Text Only(Llama-3)
Macro Precision 0.7421 $\pm$0.0047 0.7477 $\pm$0.0150 0.7136
Macro Recall 0.7433 $\pm$0.0011 0.6831 $\pm$0.0216 0.6563
Macro F1 0.7396 $\pm$ 0.0016 0.7112 $\pm$0.0095 0.6739
Micro Precision 0.7550 $\pm$0.0026 0.7650 $\pm$ 0.0201 0.7145
Micro Recall 0.7809 $\pm$0.0027 0.7372 $\pm$0.0195 0.7175
Micro F1 0.7678 $\pm$0.0026 0.7412 $\pm$ 0.0209 0.7160
Table 3: Overall performance comparison across different models.
@@
For Anger, Text Only (Llama-3) achieved the highest precision (0.7337), but Text + Explanation (RoBERTa) outperformed in recall (0.6304) and F1-score (0.6479), demonstrating better overall performance. Text Only (RoBERTa) had a lower recall (0.5116) and F1-score (0.5871), while Llama-3 performed the worst with significantly lower recall (0.4193) and F1-score (0.5336), likely due to its difficulty in capturing the nuances of Anger and the smaller dataset available.
For Fear, Text Only (RoBERTa) achieved the highest precision (0.8009), while Text + Explanation (RoBERTa) excelled in recall (0.8739) and F1-score (0.8343), demonstrating the best overall performance. Llama-3 underperformed across all metrics, with lower precision, recall, and F1-score. The larger dataset for Fear helped Text + Explanation (RoBERTa) perform well, while Llama-3 struggled to effectively utilize this data.
@@
For Joy, Text + Explanation (RoBERTa) led in recall (0.7291) and F1-score (0.7581), while Text Only (Llama-3) had the highest precision (0.7971). Despite Llama-3 excelling in precision, its lower recall (0.6567) and F1-score (0.7201) showed it lagged behind RoBERTa. The distinct nature of Joy likely aided both models’ performance.
For Sadness, Text Only (RoBERTa) had the best precision (0.7636), while Text + Explanation (RoBERTa) excelled in recall (0.8127) and F1-score (0.7423). Llama-3 performed worse across all metrics, with precision (0.6624), recall (0.7662), and F1-score (0.7105) lower than both RoBERTa models. The higher performance of Text + Explanation (RoBERTa) was likely due to its richer context, which helped capture the nuances of Sadness better than Llama-3.
In Surprise, Text + Explanation (RoBERTa) achieved the highest precision (0.7625), while Text Only (RoBERTa) outperformed in recall (0.6877). Llama-3 showed weaker performance across all metrics, with precision (0.6091), recall (0.6008), and F1-score (0.6049) significantly lower. This could be due to Surprise being less distinct than emotions like Joy, and Llama-3 struggling to capture the nuances of this emotion as effectively.
5.2 Overall Performance Comparison Across Different Models
Table 3 compares the performance of three methods: Text + Explanation (RoBERTa), Text Only (RoBERTa), and Text Only (Llama-3) across macro and micro precision, recall, and F1 scores.
Text + Explanation (RoBERTa) outperformed the other models in both Macro Recall (0.7433) and Macro F1 (0.7396), as well as in Micro Recall (0.7809) and Micro F1 (0.7678), showing its effectiveness in capturing a wide range of emotions. However, Text Only (RoBERTa) achieved the highest Macro Precision (0.7477) and Micro Precision (0.7650), indicating better selectivity in predictions. Llama-3 consistently performed worse across all metrics, particularly in recall and F1 scores.
Lotus performed well overall, with Text + Explanation (RoBERTa) excelling in emotions like Fear (F1: 0.8343), Joy (F1: 0.7581), and Sadness (F1: 0.7423), reflecting a strong ability to balance precision and recall. Even in Anger, where limited data affected recall, Text + Explanation (RoBERTa) outperformed other models with an F1-score of 0.6479. While Llama-3 performed reasonably well for Joy, it consistently lagged behind RoBERTa-based models in both precision and recall, particularly for more nuanced emotions. These results emphasize the power of combining Llama-3’s generative explanations with RoBERTa’s established strength in emotion detection. The additional layer of explanation enhances RoBERTa’s ability to capture subtle emotional nuances, improving overall performance.
In our third experiment, we fine-tuned Llama-3 for the task and compared its performance to using Llama-3’s generated explanations in conjunction with RoBERTa. Interestingly, when we used Llama-3’s explanations as input to RoBERTa, the performance exceeded that of fine-tuning Llama-3 alone. This demonstrates that combining Llama-3’s ability to generate contextual explanations with RoBERTa’s emotion detection capabilities creates a more effective approach for emotion classification tasks.
Future improvements could focus on better handling underrepresented emotions like Anger and refining the process of generating and leveraging explanations. By combining Llama-3’s explanatory power with RoBERTa’s emotion detection, we can develop a more robust and accurate solution for emotion classification tasks.
Since Anger and Surprise were the most challenging emotions for all methods, we performed error analysis for these two emotions, and the results can be found in the Appendix B.
@@ 6 Conclusion and Future Work
We proposed a novel approach for multi-label emotion detection, using Llama-3 to generate explanatory context, which was then combined with RoBERTa for emotion classification. Our method significantly improves model performance, particularly in handling ambiguous emotional expressions and multi-label tasks, outperforming text-only models in key metrics like F1-scores for fear, joy, and sadness.
Future work will focus on improving the detection of underrepresented emotions like anger and refining the explanation generation process to capture subtler emotional cues. Additionally, exploring techniques for handling imbalanced datasets and expanding the approach to different languages and emotional contexts will enhance the model’s generalizability and robustness across diverse applications.

### v3

@@ 3.2 Phase 2: RoBERTa Fine-Tuning for Multi-Label Emotion Classification
Method Macro Micro
Precision Recall F1 Precision Recall F1
Text + Exp (LLaMA-3) + RoBERTa 0.7421 $\pm$ 0.0047 0.7433 $\pm$ 0.0011 0.7396 $\pm$ 0.0016 0.7550 $\pm$ 0.0026 0.7809 $\pm$ 0.0027 0.7678 $\pm$ 0.0026
Text Only (RoBERTa) 0.7477 $\pm$ 0.0150 0.6831 $\pm$ 0.0216 0.7112 $\pm$ 0.0095 0.7650 $\pm$ 0.0201 0.7372 $\pm$ 0.0195 0.7412 $\pm$ 0.0209
Text Only (LLaMA-3) 0.7136 0.6563 0.6739 0.7145 0.7175 0.7160
Text + Exp (Mistral) + RoBERTa 0.7719 $\pm$ 0.0051 0.7206 $\pm$ 0.0135 0.7436 $\pm$ 0.0068 0.7889 $\pm$ 0.0028 0.7608 $\pm$ 0.0083 0.7746 $\pm$ 0.0037
Table 2: Overall performance comparison across different models.
@@ 4 Experimental Setup
We evaluated our multi-label emotion detection approach by fine-tuning RoBERTa with explanatory content generated by LLaMA-3 on the BRIGHTER dataset.
Text preprocessing and tokenization were performed with the RobertaTokenizer from Hugging Face. In the first phase, LLaMA-3 generated explanations, which were concatenated with the original text. In the second phase, both the original text and the generated explanations were tokenized together, allowing the model to learn the emotional context.
For fine-tuning LLaMA-3, we used 4-bit quantization and LoRA, with a batch size of 2, gradient accumulation steps of 4, and a learning rate of $1\times 10^{-4}$ for 30 training steps. These explanations were then used in the second phase for emotion classification. RoBERTa was fine-tuned with binary emotion labels (0 or 1) for each emotion in the dataset, using a batch size of 8, a learning rate of $5\times 10^{-5}$, and 3 epochs. The model performance was evaluated using precision, recall, and F1-scores, including both Macro and Micro F1-scores to assess multi-label classification.
@@
Emotion Text + Exp (LLaMA-3) + RoBERTa Text Only (RoBERTa) Text Only (LLaMA-3) Text + Exp (Mistral) + RoBERTa
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Anger 0.6695 0.6304 0.6479 0.6892 0.5116 0.5871 0.7337 0.4193 0.5336 0.7196 0.6056 0.6577
Fear 0.7983 0.8739 0.8343 0.8009 0.8200 0.8149 0.7658 0.8387 0.8006 0.8238 0.8601 0.8416
Joy 0.7957 0.7291 0.7581 0.7587 0.6925 0.7232 0.7971 0.6567 0.7201 0.7687 0.7687 0.7687
Sadness 0.6831 0.8127 0.7423 0.7636 0.6935 0.7268 0.6624 0.7662 0.7105 0.7743 0.7321 0.7526
Surprise 0.7625 0.6702 0.7132 0.7248 0.6877 0.7039 0.6091 0.6008 0.6049 0.7378 0.6834 0.7096
@@ 5 Results
In this section, we present the performance of our system, Lotus, on the competition task. Using the Text + Explanation (RoBERTa) method, Lotus achieved a score of 0.7319, outperforming the SemEval Baseline (0.7083), but falling short of the top score of 0.823. Ranked 36th, Lotus performed competitively, although there is still room for improvement to reach the top positions.
5.1 Overall Performance Comparison Across Models
Table 2 provides a summary of the overall performance of Lotus across four methods:
Text + Explanation (LLaMA-3) + RoBERTa: In this approach, LLaMA-3 generates explanations, and these explanations are combined with the original text to fine-tune RoBERTa for emotion classification.
Text Only (RoBERTa): This model uses only the text (without any explanations) to fine-tune RoBERTa.
Text Only (LLaMA-3): This model fine-tunes LLaMA-3 directly with text for emotion classification.
Text + Explanation (Mistral) + RoBERTa: In this method, Mistral generates explanations, which are combined with the original text and used to fine-tune RoBERTa.
Among these methods, Text + Explanation (LLaMA-3) + RoBERTa achieved the best overall performance, with Macro F1 (0.7396) and Micro F1 (0.7678). This approach outperformed the other methods in both recall and F1-score, demonstrating the value of combining LLaMA-3’s generative explanations with RoBERTa’s emotion detection capabilities.
Text Only (RoBERTa) achieved the highest Macro Precision (0.7477) and Micro Precision (0.7650), indicating better selectivity in its predictions. However, it lagged behind in recall and F1-scores, particularly when compared to Text + Explanation (LLaMA-3) + RoBERTa.
Text Only (LLaMA-3) performed the weakest overall, especially in recall and F1-scores. This highlights the limitations of fine-tuning LLaMA-3 directly with text without the added benefit of explanations.
Text + Explanation (Mistral) + RoBERTa showed performance similar to Text + Explanation (LLaMA-3) + RoBERTa, with slight improvements in recall and F1-score. However, the difference between Mistral and LLaMA-3 was minimal and may not be significant, suggesting that both models can perform similarly when combined with RoBERTa for fine-tuning.
@@ 5.2 Performance Comparison for Individual Emotions
Table 3 compares performance across individual emotions, showing precision, recall, and F1-scores for each method.
@@
Anger: Text + Explanation (LLaMA-3) + RoBERTa achieved precision (0.6695), recall (0.6304), and F1-score (0.6479). Text Only (RoBERTa) had the highest precision (0.6892), but lower recall (0.5116) and F1-score (0.5871). Text + Explanation (Mistral) + RoBERTa performed similarly to LLaMA-3, with precision (0.7196), recall (0.6056), and F1-score (0.6577).
Fear: Text + Explanation (LLaMA-3) + RoBERTa achieved the highest recall (0.8739) and F1-score (0.8343), outperforming Text Only (RoBERTa), which had lower recall (0.3200) and F1 (0.5149). Text + Explanation (Mistral) + RoBERTa showed slight improvements in recall (0.8601) and F1-score (0.8416), but the difference with LLaMA-3 was marginal.
@@
Joy: Text + Explanation (LLaMA-3) + RoBERTa led in recall (0.7291) and F1-score (0.7581), while Text Only (LLaMA-3) excelled in precision (0.7971). Despite LLaMA-3’s higher precision, it had lower recall (0.6567) and F1 (0.7201), trailing behind Text + Explanation (LLaMA-3) + RoBERTa. Text + Explanation (Mistral) + RoBERTa showed similar performance with an F1-score of 0.7687.
Sadness: Text Only (RoBERTa) had the highest precision (0.7636), while Text + Explanation (LLaMA-3) + RoBERTa excelled in recall (0.8127) and F1-score (0.7423). Text + Explanation (Mistral) + RoBERTa showed slight improvements in recall (0.7321) and F1-score (0.7526), with minimal differences compared to LLaMA-3.
Surprise: Text + Explanation (LLaMA-3) + RoBERTa achieved the highest precision (0.7625), while Text Only (RoBERTa) had the highest recall (0.6877). Text + Explanation (Mistral) + RoBERTa showed improved precision (0.7378) and recall (0.6834). LLaMA-3 performed weakest with precision (0.6091), recall (0.6008), and F1-score (0.6049), likely due to its difficulty in capturing the nuances of Surprise compared to other emotions.
6 Discussion
We introduced Lotus, a multi-label emotion detection approach combining LLaMA-3’s generative explanations with RoBERTa for emotion classification. This combination significantly improved performance, particularly for nuanced emotions like Fear (F1: 0.8343), Joy (F1: 0.7581), and Sadness (F1: 0.7423), surpassing text-only models.
Integrating LLaMA-3’s explanations with RoBERTa effectively balanced precision and recall, outperforming Text Only (LLaMA-3), especially for complex emotions like Fear and Sadness, emphasizing the importance of explanatory context in capturing emotional nuances.
Although LLaMA-3 was initially chosen, smaller models like Mistral and Qwen faced no significant GPU constraints on Kaggle. After testing Mistral, the results were nearly identical to LLaMA-3, suggesting both models perform similarly when fine-tuned with RoBERTa. Further exploration of other models will provide more insights.
For further illustration, Table 4 in the Appendix presents input sentences, predicted emotions, and generated explanations, providing context to clarify emotional intent and improve classification accuracy.
@@ 7 Conclusion and Future Work
Lotus showed that combining generative explanations with emotion detection models significantly improves emotion classification, particularly for ambiguous emotions. Using LLaMA-3 for explanation generation and RoBERTa for emotion detection enhanced the system’s ability to handle nuanced emotional expressions.
Future work will focus on improving detection of underrepresented emotions like Anger, refining the explanation generation process, and addressing imbalanced datasets. Expanding the model to support multiple languages and emotional contexts will enhance its generalizability. Additionally, we plan to compare Mistral with other models like Qwen and conduct ablation studies to assess their contributions. Further improvements will target challenging emotions like Anger and Surprise, with error analysis and model comparisons refining the system.